OpenStudy (anonymous):

If sin A = -8/17 with 180

myininaya (myininaya):

\[\sin(\frac{A}{2})=\pm \sqrt{\frac{1-\cos(A)}{2}}\] \[\cos(\frac{A}{2})=\pm \sqrt{\frac{1+\cos(A)}{2}}\] Now we have A is between 180 and 270 which means A/2 is between 180/2=90 and 270/2=135 So you should be able to determine what signs cos(A/2) and sin(A/2) will be based on that.

myininaya (myininaya):

Do you need help finding cos(A) given sin(A)=-8/17 with A between 180 and 270?

OpenStudy (anonymous):

Yes, please!

OpenStudy (anonymous):

A = 208 deg

OpenStudy (anonymous):

That is because \[\sin ^{-1}(-8/17) = 208 \deg\]

OpenStudy (anonymous):

What about cos A/2 and tan A/2

OpenStudy (anonymous):

Do you have a calculator?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

A/2 = 104 deg, so take cos(104) and tan(104)

OpenStudy (anonymous):

i get -.95 for cos?

OpenStudy (anonymous):

and .33 for tan

OpenStudy (anonymous):

Ack! You're using radians, not degrees. Set your calculator to degrees.

OpenStudy (anonymous):

and try again.

OpenStudy (anonymous):

not even sure how to go from radians to degrees

OpenStudy (anonymous):

What kind of calculator are you using?

OpenStudy (anonymous):

i have a ti-83

OpenStudy (anonymous):

O good. From the calculator main screen, press mode, then down 4, to angle; it should say RADIAN. You want it to say DEGREE.

OpenStudy (anonymous):

now i get cos-.24

OpenStudy (anonymous):

Great!

OpenStudy (anonymous):

so its -24 degrees? for cos a/2?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

didnt think so!

OpenStudy (anonymous):

-.24 is a length. It is the x component of a radian propped up to 104 degrees. I'll draw you a picture

OpenStudy (anonymous):

you are awesome thanks! have a big midterm tomorrow and stuck on a few problems

OpenStudy (anonymous):



OpenStudy (anonymous):

Sure, I'm glad to help. I got my math degree from MIT about 10 years ago, so it's nice to see current HS level math problems and help out with answers :-)

OpenStudy (anonymous):

this is college!

OpenStudy (anonymous):

Oh, srry!

OpenStudy (anonymous):

i guess i am still confused on where to plug thijngs in?

OpenStudy (anonymous):

Try using your TI to find the value for cos(90), sin(180), and tan(45). To start, make sure you are in DEGREE mode. Then press cos and 9 and 0 then return.

OpenStudy (anonymous):

for cos 90 i get 0

OpenStudy (anonymous):

Exact!

OpenStudy (anonymous):

sin180 i get o, tan 45 i get 1

OpenStudy (anonymous):

exact and exact. Now try sin(104) and tan(104). You've already done a good job getting cos(104), which is -.24

OpenStudy (anonymous):

Remember we are plugging in 104 because A = 208 and the problem asks for sin(A/2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

for sin104 i get .97

OpenStudy (anonymous):

yes .97 is right

OpenStudy (anonymous):

so sin a/2 is .97 degrees? doesnt seem right

OpenStudy (anonymous):

Um, .97 is a measurement of the length of the height of a triangle. It's not .97 degrees, it's .97. sin and cos take angles, and convert them into length measurements.

OpenStudy (anonymous):

guess i need to get a tutor

OpenStudy (anonymous):

well, I can try to answer as many questions as you have for now. I'm going to be up late anyway.

OpenStudy (anonymous):

im trying to solve this one first for sin cos and tan

OpenStudy (anonymous):

You have sin and cos already. Now try tan.

OpenStudy (anonymous):

i must be missing something bc sin a/2 given the equation above im getting decimals

OpenStudy (anonymous):

tan is -4

OpenStudy (anonymous):

sin and cos always give answers between -1 and 1. Just because they are decimals doesn't mean they are in degrees.

OpenStudy (anonymous):

The max value of cos(x) for any x is 1 The max value of sin(x) for any x is 1 The min value of cos(x) for any x is -1 and the min value of sin(x) for any x is -1

OpenStudy (anonymous):

that i know, but cant seem to find the connect w this problem